How many milliliters of 12.0 M HCl should be added to 5.00 x 102 mL of 0.100 M Na2SO3 to produce a pH 7.00 buffer solution. For H2SO3 Ka1 = 1.39 x 10-2 and Ka2 = 6.73 x 10-8.
Answer #1

Answer : We know , moles = molarity (M) x volume (v in L) So, Moles of S03- 0.100 O:500 %3D = O.050 mol. moles of Ht to be ad= log (e-05=x) 0.05-x => 7 = 7.172 -0:172 = log 0.05-X) -> O.673 O. 05-X %3D = O:05 O.67 3 X + x 0.05 1.673 mol O.OR99 molesPlease do give a thumbs up.